OVHCloud Network Status

Current status
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
Scheduled Maintenance Report for Network & Infrastructure
We noticed CRC errors between one on the two links which connect chi-5-n7 et chi-1-n7.
We are going to clean the fiber.
The second link has the capacity to handle all traffic.
Our customers will not be impacted by this maintenance.


Date: 2017-01-06 20:54:53 UTC
Despite having change all equipments in the link between the two equipments, we still see CRCs. We are going to send new equipments (2 Connectors + fiber) and replace the whole link when ready. A new task will be open when we will proceed.
Meanwhile the link has been reactivated to ensure a service redundancy.

Date: 2017-01-04 19:57:47 UTC
The optical fiber has been cleaned and CRC are still present. We are going to switch the fiber connector on the equipment chi-5-n7.
Posted Jan 04, 2017 - 16:56 UTC